RAX-2 operational status will be recorded below. We plan to provide brief updates on scheduled RAX-2 operations and current status. Some additional links proving related information include: RAX TLEs || Live Telemetry || Blog.
Date: Friday 2013/03/18, Status: Multiple experiments performed since the last update, but no FAI detected. Expt 1031 was run over the past weekend and is pending download. RAX had an unexpected reboot around the time of the experiment, likely due to a single event upset.
|2013/02/01||Expt 1026 completed with PFISR January 18, no echo detected. We are currently downloading the last portion of raw data from expt1025 while keeping and eye on space weather forecasts. Experiments will be run when the required space weather conditions arise.|
|2012/12/22||Radar data downloads were paused over the week to log and download GPS data. The testing was successful and we’re back to downloading unprocessed radar data from previous experiments over the semester break.|
|2012/12/05||Experiment 1025 was performed with MUIR on November 11; no FAI detected. We are currently downloading unprocessed radar samples from experiment 1022, in which scintillation of MUIR radar pulses was detected.|
|2012/11/05||Two experiments with RISR have been performed since the last update below. Space weather conditions were calm and no FAI were detected. We’re currently downloading radar measurements from experiment 1022 and have more experiments planned for later this month.|
|2012/09/24||The data from expt1022 (MUIR/HAARP) has been transferred to the IDPU for safekeeping, and we are finishing up the download of raw data from expt1019.|
|2012/08/17||Two experiments with MUIR were performed last week (1021 and 1022). The download is raw data from experiment 1019 is still being downloaded. When complete we will download data from 1022, the first detection of artificial FAI generated by HAARP.|
|2012/07/23||Two experiments were performed 7/14-7/15, one of them was processed and downloaded, but no echoes were detected. During all time other than these experiments, we have been downloading more expt1019 radar samples.|
|2012/06/21||We finished downloading all the expt1018 raw data yesterday and are starting the download for expt1019, which was the most recent experiment performed and detected radar scatter.|
|2012/05/17||This week we are transferring some more radar measurement data from the most recent experiment for download, which will allow us to resume experiments without over-writing valuable data. When this is complete, we will continue daily downloads of science data that remains onboard from recent radar experiments.|
|2012/04/28||Geomagnetic activity increased over the past weak, so Expt 1018 was performed with PFISR on 4/25. Radar echoes detected! The processing and download of processed data was completed within a day. We have now resumed downloading raw radar measurements.|
|2012/04/22||We’ve re-enabled scheduled downloads without the SD card and are now downloading data from expt1017 over amateur radio operators worldwide. Downloads will continue as space weather activity remains calm.|
|2012/03/20||We are still downloading more raw data from expt1017. The download will likely continue over the next few weeks.|
|2012/03/12||We are downloading some more detailed data from expt1017 for further analysis. Expt1016 was also downloaded over the weekend, and 1015 remains onboard from earlier in the week. Experiments will resume after the download of 1017 data.|
|2012/03/07||After a brief break, we’ve returned to performing experiments, and completed Experiment 1016 yesterday and plan to perform Experiment 1017 today. There are reports of interesting solar activity, so we are excited by the outcome of the experiments over the next couple days.|
|2012/02/29||Experiment 1015 ran today with PFISR, but there was minimal geomagnetic activity so the experiment will not be processed and downlinked. Space weather is forecasted to be calm through the weekend so we plan to do the next experiment next week. In the mean time, we are downloading a few pieces of experiments that are still onboard from before the SD anomaly.|
|2012/02/23||Experiment 1014 processing is complete and we should finish downloading it today. Then we will finish up downloading two sets of experiment data that are still onboard from before the SD anomaly and schedule the next experiment.|
|2012/02/20||Experiment 1014 is scheduled for today, and during the rest of the week, we will process and download 1014 along with downloading the remainder of the science data that’s still onboard from before the anomaly.|
|2012/02/16||We are performing the final tests on new software that will enable R2 to resume regular science experiments without using the SD card. We expect to schedule the next science experiment this weekend.|
|2012/01/19||R2 normal operations are temporarily on hold as we debug an error on the flight computer SD card.|
|2012/01/11||RAX-2 has moved into a phase of performing science experiments nearly every day! Over the past 7 days, we have performed 6 experiments with PFISR. Every experiment is being processed on board and is then in the queue for download when R2 passes over ground stations.|
|2012/01/06||R2 performed an experiment with PFISR today, we are currently processing the data, and another experiment is scheduled for 2012/01/07.|
|2012/01/05||RAX-2 is set for it’s next experiment 2012/01/06, and depending on the space weather conditions at that time, we will run another experiment on 2012/01/07.|
|2012/01/02||RAX has returned from a brief break. R2 performed experiment 1004 on 2011/12/29, see the RTI results here. High winds (up to 50 mph) prevented significant commanding from our Ann Arbor station on 2012/01/01. We are currently downloading beacons over our global network and performing another Short Duration Analysis Package (SDAP). An SDAP collects important satellite health information (such as attitude determination sensor data, voltages, currents, and temperatures) at 1 Hz for 20 minutes when R2 is in the sun. We plan to perform another NoFlo experiment over the Pacific ocean.|
|2011/12/19||R2 performed experiment 1003 with PFISR on Sunday 12/18. This was the third run of 1003 as the first two were during quiet space weather. For Sunday’s experiment, additional PFISR pulses were used for simultaneous ionospheric measurements. Space weather conditions were calm, but we are processing the data to check the effect of the additional pulses. The next experiment is scheduled for 2011/12/28. This week, we are testing the polarization switch for the UHF antenna and measuring the resulting antenna performance.|
|2011/12/16||R2 performed two experiments this week with PFISR, but both were during quiet space weather conditions so were not processed. We will resume experiments on 2011/12/18. In the meantime, we are performing NoFlos — noise floor tests — to characterize the noise floor of the radar receiver. This information will be used in analyzing the scientific data.|
Beacons: RAX-2 collects data from sensors throughout the spacecraft every 20 seconds and transmits the information. The data is also stored on-board so it can be downloaded at a later time.
NoFlo: Noise floor test. We take data with the radar receiver in relatively quite regions of space, such as over oceans, to characterize the spacecraft-induced noise
SDAP: Short duration analysis package. This is 1 Hz data from sensors across all subsystems, normally collected for 20 minute durations. It provide higher frequency data than the beacons and is used to assess satellite health.
SOAP: Single orbit analysis package. This is similar to an SDAP but the data is collected for 100 minutes to cover an entire orbit.
(Banner image © RAX Team)